# PCWE for FSAI – Derivation of scalar wave equations for fluid-structure-acoustics interaction of low Mach number flows

### v3

@@ 1 Introduction
Aeroacoustic analogies [1] compute noise radiation efficiently but the obtained fluctuating field converges only in steady flow regions to the acoustic field [2, 3]. As recognized by Phillips [4] and Lilley [5], the source terms responsible for mean flow-acoustics interactions should be part of a convective wave operator. Therefore, an aeroacoustic approach based on a systematic decomposition of the field properties emerged. This circumvents that sources depend on the acoustic solution and provides a rigorous definition of acoustics. Ribner [6] decomposed the fluctuating pressure in a pseudo pressure and an acoustic pressure part $p^{\prime}=p^{0}+p^{\mathrm{a}}$. Hardin and Pope [7] formulated their viscous/acoustic splitting technique expansion about the incompressible flow (EIF). The EIF formulation was modified and applied to several examples [8, 9, 10]. Ewert and Schröder [11] proposed a different technique leading to the acoustic perturbation equations (APE). Instead of filtering the flow field, the source terms of the wave equations are filtered according to the characteristic properties of the acoustic modes. The acoustic modes are obtained from LEE. Recent advances of the theory led to [12]. Hüppe [13] derived a computationally efficient reformulation of the APE-2 system and named it perturbed convective wave equation (PCWE). Since then, a number of low Mach number flow applications have been addressed by computational aeroacoustics using the PCWE model. This includes human phonation [14, 15, 16, 17, 18, 19, 20, 21, 22, 23], HVAC systems [24] and fan noise [25, 26, 27, 28, 29] using openCFS [30, 31]. The PCWE is valid for incompressible flows and near-field acoustics [32]. Recently, a convective wave equation was proposed for investigating jet-noise and installation effects [33, 34, 35].
@@ 2.2 Modified scalar variant of LPCE for FSAI
$\frac{1}{c_{0}^{2}}\,\,\frac{D^{2}\psi_{\rm a}}{Dt^{2}}+\frac{1}{\rho_{0}c_{0}^{2}}\nabla\phi_{\rm a}\cdot\nabla P_{\mathrm{i}c}-\Delta\psi_{\rm a}=-\frac{1}{\rho_{0}c_{0}^{2}}\,\frac{D\Phi_{\mathrm{p}}}{Dt}\,.$ (29)

### v4

@@ 1 Introduction
Aeroacoustic analogies [1] compute noise radiation efficiently but the obtained fluctuating field converges only in steady flow regions to the acoustic field [2, 3]. As recognized by Phillips [4] and Lilley [5], the source terms responsible for mean flow-acoustics interactions should be part of a convective wave operator. Therefore, an aeroacoustic approach based on a systematic decomposition of the field properties emerged. This circumvents that sources depend on the acoustic solution and provides a rigorous definition of acoustics. Ribner [6] decomposed the fluctuating pressure in a pseudo pressure and an acoustic pressure part $p^{\prime}=p^{0}+p^{\mathrm{a}}$. Hardin and Pope [7] formulated their viscous/acoustic splitting technique expansion about the incompressible flow (EIF). The EIF formulation was modified and applied to several examples [8, 9, 10]. Ewert and Schröder [11] proposed a different technique leading to the acoustic perturbation equations (APE). Instead of filtering the flow field, the source terms of the wave equations are filtered according to the characteristic properties of the acoustic modes. The acoustic modes are obtained from LEE. Recent advances of the theory led to [12]. Hüppe [13] derived a computationally efficient reformulation of the APE-2 system and named it perturbed convective wave equation (PCWE). Since then, a number of low Mach number flow applications have been addressed by computational aeroacoustics using the PCWE model. This includes human phonation [14, 15, 16, 17, 18, 19, 20, 21, 22, 23], HVAC systems [24] and fan noise [25, 26, 27, 28, 29] using openCFS [30, 31]. The PCWE is valid for incompressible flows and near-field acoustics [32]. Recently, a convective wave equation was proposed for investigating jet-noise and installation effects [33, 34, 35, 36, 37].
@@ 2.2 Modified scalar variant of LPCE for FSAI
$\frac{1}{c_{0}^{2}}\,\,\frac{D^{2}\phi_{\rm a}}{Dt^{2}}+\frac{1}{\rho_{0}c_{0}^{2}}\nabla\phi_{\rm a}\cdot\nabla P_{\mathrm{i}c}-\Delta\phi_{\rm a}=-\frac{1}{\rho_{0}c_{0}^{2}}\,\frac{D\Phi_{\mathrm{p}}}{Dt}\,.$ (29)